OpenStudy (anonymous):

In an 45-45-90 triangle, I am giving the number 6, the letter m,and the letter n.... It says solve for th varibles....How?

OpenStudy (jdoe0001):

got any pictures to share with us?

OpenStudy (anonymous):

No I don't.

OpenStudy (jdoe0001):

in the 45-45-90 rule, all 3 sides correlate to each other by some ratio, so once you have only 1 side alone, you can find the rest really http://www.moomoomath.com/45_45_90_moomoo.JPG

OpenStudy (jdoe0001):

what does 6 stand for? the hypotenuse? any of the other 2 sides, something tells me you do have a picture :S

OpenStudy (anonymous):

6 is the hypotenuse, m is one leg, and n is the other. It says to solve for m and n.

OpenStudy (jdoe0001):

look at the picture, notice the ratios :), look closely at the ratios, the hypotenuses is "something", and the other 2 sides are equal to each other

OpenStudy (jdoe0001):

so once you find one of the sides, you find both

OpenStudy (anonymous):

This is so confusing... I honestly don't understand how to get both sides with only the number 6

OpenStudy (jdoe0001):

did you see the picture for the 45-45-90 rule? what's the hypotenuse?

OpenStudy (anonymous):

6 is the hypotenuse

OpenStudy (jdoe0001):

ok, and the one in the picture? http://www.moomoomath.com/45_45_90_moomoo.JPG

OpenStudy (anonymous):

x square root 2

OpenStudy (anonymous):



OpenStudy (jdoe0001):

so we can say from the correspondence that \(\bf 6 = x\sqrt{3}\) so what's "x"?

OpenStudy (jdoe0001):

woops wrong rule there darn

OpenStudy (anonymous):

since both angles are the same (they are both 45) then both sides are the same lets call them both \(m\) then by pythagoras \[m^2+m^2=6^2\] \[2m^2=36\] \[m^2=18\] \[m=\sqrt{18}=3\sqrt{2}\]

OpenStudy (jdoe0001):

so we can say from the correspondence that \(\bf 6 = x\sqrt{2}\)

OpenStudy (anonymous):

Im sorry but I still don't understand

OpenStudy (jdoe0001):

your hypotenuse is 6